If x, y are real numbers such that ordered pairs (x + y, x -y) and (2x + 3y, 3x - 2y) are equal, then (x, y) is equal to
1
Expert's answer
2020-04-20T11:50:38-0400

Given, the ordered pairs "(x + y, x - y)" and "(2x + 3y, 3x - 2y)" are equal. Since, two ordered pairs are equal if and only if corresponding coordinates are equal, we get "x+y=2x+3y" and "x-y=3x-2y" . That is "-x-2y = 0" and "-2x + y = 0". Solving the two equations, we get "x = 0" and "y = 0" . Hence the ordered pair "(x,y) = (0,0)".